OpenStudy (anonymous):

I WILL MEDAL Kim is solving the following linear equation. 11+3x-7=6+5-3x Her final two steps are: 4+3x=3x+5 4=5 Select the statement that correctly interprets Kim's solution. The solution is x = 0. The solution is the ordered pair (4, 5). There is no solution since 4 = 5 is a false statement. There are infinitely many solutions since 4 = 5 is a false statement

OpenStudy (calculusxy):

Is it possible to say that 4 is the same thing as 5?

OpenStudy (anonymous):

11 + 3 x - 7 = 6 + 5 - 3 x x = 7/6

OpenStudy (calculusxy):

11 + 3x - 7 = 6 + 5 - 3x 4 + 3x = 11 - 3x 4 + 6x = 11 6x = 7 x = 7/6 @robtobey says the correct solution. It doesn't make sense how Kim would get 4 = 5 as the solution to this linear equation.